# Thorough Analysis of $\pi^{-}p\rightarrow K^{*}\Sigma$ Reaction: Probing Nucleon Excited State and Constituent Scaling Laws

### v1

@@ Abstract
In this study, applying the effective Lagrangian approach, we investigate the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ to explore the production of the nucleon excited state $N^{*}(2080)$ and the constituent field scaling laws, also known as the constituent counting rules. Specifically, we analyze two scenarios, where the $J^{P}$ quantum numbers of $N^{*}(2080)$ are $3/2^{-}$ and $1/2^{-}$. Our results show that the contribution from $s$-channel $N^{*}$ exchange dominates the total cross section, surpassing those from $t$-channel $K^{(*)}$ exchange and $u$-channel $\Sigma$ exchange. Additionally, we present predictions for the differential cross section of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process. Notably, in the large momentum transfer region ($\theta=90^{\circ}$), the differential cross section in the $t$-distribution aligns with the energy-dependent scaling formula predicted by the constituent counting rule, specifically, $d\sigma/dt\sim 1/s^{8}$ for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction. Finally, we extend the 2-to-2 scattering process to the 2-to-3 Dalitz process and provide theoretical predictions for this scenario. Our findings suggest that the cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ process can reach several tens of $\mu b$ near the threshold, making it highly favorable for experimental measurement. However, the current scarcity of experimental data for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at threshold energy limits our ability to precisely determine the properties of the nucleon excited state $N^{*}$ and related physical quantities. Therefore, additional experimental measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at threshold are strongly encouraged, and correlation studies could be pursued at facilities such as J-PARC, AMBER, and upcoming HIKE and HIAF meson beam experiments.
@@ I INTRODUCTION
Over the past several decades, meson-nucleon scattering experiments have played a crucial role in the development of hadron spectroscopy wl:2024ext . These experiments are particularly significant, as more than $80\%$ of the light hadrons cataloged by the Particle Data Group (PDG) ParticleDataGroup:2022pth have been discovered or characterized through such studies. Hadron spectroscopy provides an effective means to enhance our understanding of the nonperturbative aspects of strong interactions. In recent years, theorists have also focused on scattering processes involving $K$ and $\pi$ mesons, aiming to explore the production of new hadronic states Liu:2019zoy ; Chen:2016qju ; Chen:2022asf ; Zou:2016bxw ; Lin:2022eau ; Wang:2023lia ; Wang:2022sib ; Xiang:2020phx ; Wang:2019zaw ; Wang:2019uwk ; Wang:2019dsi ; Wang:2018mjz ; Wang:2017qcw ; Cheng:2016ddp ; Wang:2015xwa ; Liu:2021ojf ; Gao:2012zh ; Wu:2009tu ; Xie:2014kja ; Xie:2007qt . Both experimental and theoretical studies suggest that meson beam experiments offer unique advantages in advancing research on hadron spectroscopy and related physical phenomena.
@@
Light-flavor baryons have garnered significant attention from the hadron physics community due to the well-known “missing” problem associated with their excited states ZOU:2005 . Among the numerous observed light-flavor baryon states, the $N^{*}(2080)$ stands out as a particularly intriguing research target, closely linked to the investigation of hidden strange pentaquark states. The nucleon excited state $N^{*}(2080)$ has been proposed as a $K^{*}\Sigma$ molecular state, potentially serving as the strange partner of the $P_{c}^{+}(4457)$ hadronic molecular state He:2017aps ; Lin:2018kcc ; Khemchandani:2011et ; Doring:2008sv . The potential existence of such hadronic molecules has also been suggested in studies utilizing the quark delocalization color screening model Gao:2017hya . As discussed in Ref. wl:2024ext , hidden strange pentaquark states could be identified and studied by measuring processes such as $\pi^{-}p\rightarrow K^{(*)}\Sigma^{(*)}$ and $\pi^{-}p\rightarrow K^{(*)}\Lambda^{(*)}$. These measurements could provide crucial experimental data to address the “missing” problem of baryon excited states. In Ref. Lin:2018kcc , the effective Lagrangian method was employed to calculate the decay modes of the $N^{*}(2080)$ having $J^{P}=3/2^{-}$ as an $S$-wave $K^{*}\Sigma$ molecular state. The results showed that the measured decay properties of $N^{*}(2080)$ with $J^{P}=3/2^{-}$ were well reproduced, supporting the molecular interpretation of this state.
Given that the $J^{P}$ quantum numbers of the $N^{*}(2080)$ [^1] According to the latest PDG ParticleDataGroup:2022pth , the previously identified two-star nucleon resonance $N^{*}(2080)$ has been split into a three-star $N(1875)$ and a two-star $N(2120)$, both with spin-parity $3/2^{-}$. A key characteristic of hadronic molecules is that they are typically very shallowly bound Guo:2017jvc , meaning that their masses are generally just below the corresponding threshold. In this study, we consider the $K^{*}\Sigma$ molecular interpretation, with a threshold of 2086 MeV, and accordingly, we adopt the mass of the $N(2120)$ as 2080 MeV, referring to it by its previous designation, the $N^{*}(2080)$, throughout the remainder of this paper. could be $1/2^{-}$ or $3/2^{-}$, this state can couple to $K^{*}\Sigma$ through $S$ wave interaction. The coupling constant between $N^{*}$ and $K^{*}\Sigma$ can be determined using the Weinberg compositeness criterion Lin:2018kcc ; Weinberg:1965zz ; Baru:2003qq ; Lin:2017mtz . Therefore, searching for and measuring the $N^{*}(2080)$ state via the $\pi^{-}p\rightarrow K^{*}\Sigma$ process is a feasible approach. In the 1970s and 1980s, several experiments measured scattering cross section data for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process Crennell:1972km ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; Dahl:1967pg ; Miller:1965 ; Abramovich:1972rq , providing crucial support for our research.
@@
In this work, we will perform calculations and analysis of the $\pi^{-}p\rightarrow K^{*}\Sigma$ scattering process using the effective Lagrangian approach, with the aim of investigating the production of the nucleon excited state $N^{*}(2080)$ and exploring the constituent counting rule. Additionally, we will analyze the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ Dalitz process and assess its experimental feasibility, thereby offering valuable theoretical insights for future experimental studies.
Actually, the $\pi^{-}p\rightarrow K^{*}\Sigma$ process serves as an ideal case for examining the constituent field scaling law, commonly known as the constituent counting rule. Initially proposed by Brodsky in 1973 and further developed in the decade following the advent of perturbative QCD (pQCD), this rule embodies the principles of conformality and scale invariance in QCD at high energies Lepage:1980fj ; Mueller:1981sg ; Lepage:1979zb ; Brodsky:1974vy ; Brodsky:1973kr ; Matveev:1973ra ; Guo:2016fqg . The constituent counting rule has been nonperturbatively derived in contexts like AdS/QCD, showcasing its broad applicability to various field theories Polchinski:2001tt . In essence, the rule predicts that in exclusive processes such as $a+b\rightarrow c+d$ involving large momentum transfers, quarks must share significant momenta to coalesce into hadrons by exchanging hard gluons. Consequently, the cross-section for such an exclusive reaction scales according to $d\sigma/dt\sim s^{2-n}f(\theta_{\mathrm{c.m.}})$, where $n=n_{a}+n_{b}+n_{c}+n_{d}$ represents the total number of constituents in particles $a$, $b$, $c$, and $d$. Here, $s$ and $t$ are the Mandelstam variables, and $\theta_{\mathrm{c.m.}}$ is the scattering angle in the center-of-mass frame. The constituent counting rule has been validated by two-body hadronic exclusive reactions conducted at the Brookhaven National Laboratory’s Alternating Gradient Synchrotron (BNL-AGS) Baller:1988tj . More broadly, this rule effectively describes the differential cross-section behavior in various hadron-hadron scattering processes within the large momentum transfer regime JeffersonLabHallA:2002vjd ; JeffersonLabHallA:2004gxp ; White:1994tj ; Amaryan:2021cnj . Applying the constituent counting rule in our analysis of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process not only deepens our understanding of the mechanisms underlying hadron-hadron scattering and the role of constituent quarks but also allows us to assess the accuracy of our theoretical cross-section predictions.
@@
This paper is organized as follows. After the introduction, the Lagrangians and amplitudes used in this work are presented in Section II. The numerical results for the total and differential cross sections are provided in Section III. In Section IV, we introduce the constituent counting rule and assess its validity. The Dalitz process and its experimental feasibility are discussed in Section V, followed by a brief summary in Section VI.
II THE PRODUCTION OF THE $N^{*}(2080)$ VIA THE $\pi^{-}p\rightarrow K^{*}\Sigma$ REACTION
@@
The schematic tree-level Feynman diagrams for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction are depicted in Fig. 1. These include $s$-channel $N^{*}(2080)$ ($\equiv N^{*}$) exchange, $u$-channel $\Sigma$ exchange, and $t$-channel $K$ and $K^{*}$ exchanges.
@@
To evaluate the pion-induced production of $K^{*}\Sigma$, the relevant Lagrangians are required, as used in previous studies Xiang:2020phx ; Kim:2015ita ; He:2017aps ; Xie:2014kja ; Ben:2023uev ; Wang:2019dsi ; Cheng:2016ddp . In general, similar to the case of the $P_{c}$ pentaquark state, the $J^{P}$ quantum numbers of the $N^{*}(2080)$ state can be either $\frac{3}{2}^{-}$ or $\frac{1}{2}^{-}$. Therefore, for the $s$-channel, both quantum number assignments for the $N^{*}(2080)$ should be considered in our calculations. These Lagrangians include
$\mathcal{L}^{3/2^{-}}_{K^{*}\Sigma N^{*}}$ $=$ $g_{K^{*}\Sigma N^{*}}^{3/2^{-}}\bar{N}^{*}_{\mu}{\Sigma}K^{*\mu}+\mathrm{H.c.},$ (1)
$\mathcal{L}^{3/2^{-}}_{\pi NN^{*}}$ $=$ $\frac{g_{\pi NN^{*}}^{3/2^{-}}}{m_{\pi}}\bar{N}\gamma_{5}\bm{\tau}\cdot\partial^{\mu}\bm{\pi}N^{*}_{\mu}+\mathrm{H.c.},$ (2)
$\mathcal{L}^{1/2^{-}}_{K^{*}\Sigma N^{*}}$ $=$ $g_{K^{*}\Sigma N^{*}}^{1/2^{-}}\bar{N}^{*}\gamma_{5}\gamma^{\mu}K^{*}_{\mu}\Sigma+\mathrm{H.c.},$ (3)
$\mathcal{L}^{1/2^{-}}_{\pi NN^{*}}$ $=$ $g_{\pi NN^{*}}^{1/2^{-}}\bar{N}\bm{\tau}\cdot\bm{\pi}N^{*}+\mathrm{H.c.}.$ (4)
@@
where $\bm{\tau}$ denotes the Pauli matrices, and the matrix $\bm{T}$ is defined as in Ref. Matsuyama:2006rp . The symbols $\bm{\pi}$, $K$, and $K^{*}$ represent the $\pi$, $K$, and $K^{*}$ meson fields, respectively. Similarly, $N$, $\Sigma$, and $N^{*}$ correspond to the nucleon, $\Sigma$ baryon, and $N^{*}(2080)$ fields, respectively.
Assuming the $N^{*}(2080)(\frac{3}{2}^{-})$ is a pure $S$-wave molecular state of $K^{*}$ and $\Sigma$, the coupling constant $g_{K^{*}\Sigma N^{*}}^{3/2^{-}}$ can be estimated model-independently using the Weinberg compositeness criterion, yielding Weinberg:1965zz ; Baru:2003qq ; Lin:2017mtz ; Lin:2018kcc
$g_{K^{*}\Sigma N^{*}}^{2}$ $=$ $\frac{4\pi}{4M_{N^{*}}M_{\Sigma}}\frac{(M_{K^{*}}+M_{\Sigma})^{5/2}}{(M_{K^{*}}M_{\Sigma})^{1/2}}\sqrt{32\epsilon},$ (11)
$\epsilon$$\equiv$ $M_{K^{*}}+M_{\Sigma}-M_{N^{*}},$ (12)
where $M_{N^{*}}$, $M_{K^{*}}$, and $M_{\Sigma}$ represent the masses of the $N^{*}(2080)(\frac{3}{2}^{-})$, $K^{*}$, and $\Sigma$, respectively, while $\epsilon$ denotes the $K^{*}\Sigma$ binding energy. Assuming the physical state is a pure $S$-wave hadronic molecule, the relative uncertainty of the above approximation for the coupling constant is given by $\sqrt{2\mu\epsilon}\cdot r$, where $\mu=\frac{m_{1}m_{2}}{m_{1}+m_{2}}$ is the reduced mass of the bound particles, and $r$ is the range of forces, which can be estimated as the inverse of the mass of the exchanged particle. For the $K^{*}\Sigma$ system, $r$ can be estimated as $1/m_{K}$ Lin:2018kcc .
Following Ref. Lin:2018kcc , we adopt the mass of the $N^{*}(2080)(\frac{3}{2}^{-})$ as $M_{N^{*}}=2080\text{ MeV}$. Substituting this into Eq. (11), one obtains:
$g_{K^{*}\Sigma N^{*}}^{3/2^{-}}=1.72.$ (13)
Since the $N^{*}(2080)(\frac{1}{2}^{-})$ can also be considered a $S$-wave molecular state of $K^{*}$ and $\Sigma$, we approximate the value of $g_{K^{*}\Sigma N^{*}}^{1/2^{-}}$ as 1.72. In the absence of experimental measurements for the partial width of the $N^{*}(2080)$ decays to $\pi N$, we treat the coupling constants $g_{\pi NN^{*}}^{3/2^{-}}$ and $g_{\pi NN^{*}}^{1/2^{-}}$ as free parameters to be determined through fitting.
@@
$F_{t/u}(q)=\frac{\Lambda{{}^{2}_{t/u}}-m^{2}}{\Lambda{{}^{2}_{t/u}}-q^{2}}.$ (16)
@@
$\mathcal{A}^{N^{*({3/2}^{-})}}_{s,\mu}$ $=$ $-i\sqrt{2}g_{K^{*}\Sigma{N^{*}}}^{{3/2}^{-}}\frac{g_{\pi N{N^{*}}}^{{3/2}^{-}}}{m_{\pi}}F_{s}(q)\gamma_{5}k_{1}^{\nu}$ (19)
×$\times\frac{(q\mkern-8.0mu/_{s}+m_{N^{*}})}{s-m^{2}_{N^{*}}+im_{N^{*}}\Gamma_{N^{*}}}\Delta_{\mu\nu},$
$\mathcal{A}^{N^{*({1/2}^{-})}}_{s,\mu}$ $=$ $-\sqrt{2}g_{K^{*}\Sigma{N^{*}}}^{{1/2}^{-}}g_{\pi N{N^{*}}}^{{1/2}^{-}}F_{s}(q)\gamma_{5}\gamma_{\mu}$ (20)
×$\times\frac{(q\mkern-8.0mu/_{s}+m_{N^{*}})}{s-m^{2}_{N^{*}}+im_{N^{*}}\Gamma_{N^{*}}},$
$\mathcal{A}^{\Sigma}_{u,\mu}$ $=$ $-\sqrt{2}\frac{g_{\Sigma\Sigma\pi}}{m_{\pi}}g_{K^{*}N\Sigma}F_{u}(q)\gamma^{5}\gamma^{\nu}k_{1\nu}\frac{(q\mkern-8.0mu/_{\Sigma}+m_{\Sigma})}{{u-m_{\Sigma}^{2}}}$ (21)
@@
$\mathcal{A}^{K}_{t,\mu}$ $=$ $-\sqrt{2}g_{K^{*}K\pi}g_{K\Sigma N}F_{t}(q)\frac{1}{{t-m_{K}^{2}}}\gamma_{5}$ (22)
@@
$\mathcal{A}^{K^{*}}_{t,\mu}$ $=$ $-i\sqrt{2}g_{\pi K^{*}K^{*}}g_{K^{*}N\Sigma}F_{t}(q)(\gamma_{\xi}-\frac{k_{K^{*}{\Sigma}N}}{2m_{N}}\gamma_{\xi}q\mkern-8.0mu/_{K^{*}})$ (23)

### v2

# Production potential of hidden-strange molecular pentaquarks through the $\pi^{-}p\rightarrow K^{*}\Sigma$ process

@@ Abstract
In this study, applying the effective Lagrangian approach, we investigate the reaction $\pi^{-}p\rightarrow K^{*}\Sigma$ to explore the production of the hidden-strange molecular pentaquarks $P_{s\bar{s}}$. Specifically, we analyze two scenarios, where the $J^{P}$ quantum numbers of $P_{s\bar{s}}$ are $3/2^{-}$ and $1/2^{-}$. Our results show that the contribution from $s$-channel $P_{s\bar{s}}$ exchange dominates the total cross section, surpassing those from $t$-channel $K^{(*)}$ exchange and $u$-channel $\Sigma$ exchange. Additionally, we present predictions for the differential cross section of the $\pi^{-}p\rightarrow K^{*}\Sigma$ process. Finally, we extend the 2-to-2 scattering process to the 2-to-3 Dalitz process and provide theoretical predictions for this scenario. Our findings suggest that the cross section for the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ process can reach several tens of $\mu b$ near the threshold, making it highly favorable for experimental measurement. However, the current scarcity of experimental data for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at threshold energy limits our ability to determine the properties of the hidden-strange molecular pentaquarks $P_{s\bar{s}}$ and related physical quantities. Therefore, additional experimental measurements of the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction at threshold are strongly encouraged, and correlation studies could be pursued at facilities such as J-PARC, AMBER, and upcoming HIKE and HIAF meson beam experiments.
@@ I INTRODUCTION
Hadron spectroscopy offers an effective approach to deepen our understanding of the nonperturbative aspects of strong interactions. Over the past several decades, meson-nucleon scattering experiments have played a crucial role in the advancement of hadron spectroscopy wl:2024ext . This is evident from the fact that over $80\%$ of the light hadrons cataloged by the Particle Data Group (PDG) ParticleDataGroup:2022pth have been discovered through such scattering processes. While collider experiments are the primary means of exploring the nature of matter, scattering experiments deserve greater attention. In recent years, theorists have increasingly focused on scattering processes involving $K$ and $\pi$ meson beams, aiming to investigate the production of new hadronic states Liu:2019zoy ; Chen:2016qju ; Chen:2022asf ; Zou:2016bxw ; Lin:2022eau ; Wang:2023lia ; Wang:2022sib ; Xiang:2020phx ; Wang:2019zaw ; Wang:2019uwk ; Wang:2019dsi ; Wang:2018mjz ; Wang:2017qcw ; Cheng:2016ddp ; Wang:2015xwa ; Liu:2021ojf ; Gao:2012zh ; Wu:2009tu ; Xie:2014kja ; Xie:2007qt .
With the observation of several $P_{c}$ states by the LHCb Collaboration LHCb:2015yax ; LHCb:2019kea , hidden-charm molecular-type pentaquarks, predicted in Refs. Chen:2015loa ; Chen:2015moa ; Wang:2019krd ; Shimizu:2016rrd ; Roca:2015dva ; Wang:2019dsi , have been identified. This identification stems from the discovery of a characteristic mass spectrum in the $J/\psi p$ invariant mass distribution from the decay $\Lambda_{b}\to J/\psi pK$ LHCb:2015yax ; LHCb:2019kea . This represents an important progress in the study of hadron spectroscopy.
@@
Following this, it is natural to further investigate the hidden-strange partners of these hidden-charm molecular-type pentaquarks while developing the “Particle Zoo 2.0” version. Among the numerous observed light-flavor baryon states, the $N^{*}(2080)$ stands out as a particularly intriguing research target, closely linked to the investigation of hidden strange pentaquark states. The nucleon excited state $N^{*}(2080)$ has been proposed as a $K^{*}\Sigma$ molecular state, potentially serving as the strange partner of the $P_{c}^{+}(4457)$ hadronic molecular state He:2017aps ; Lin:2018kcc ; Khemchandani:2011et ; Doring:2008sv . The potential existence of such hadronic molecules has also been suggested in studies utilizing the quark delocalization color screening model Gao:2017hya . In Ref. Lin:2018kcc , the effective Lagrangian method was employed to calculate the decay modes of the $N^{*}(2080)$ having $J^{P}=3/2^{-}$ as an $S$-wave $K^{*}\Sigma$ molecular state. The results showed that the measured decay properties of $N^{*}(2080)$ with $J^{P}=3/2^{-}$ were well reproduced, supporting the molecular interpretation of this state. We observe a change in the resonance designation relation to $N^{*}(2080)$. According to the latest PDG ParticleDataGroup:2022pth , the previously identified two-star nucleon resonance $N^{*}(2080)$ has been split into a three-star $N(1875)$ and a two-star $N(2120)$, both with spin-parity $3/2^{-}$. Therefore, in our analysis, we no longer use $N^{*}(2080)$ to represent the hidden-strange molecular-type pentaquarks as considered in Refs. He:2017aps ; Lin:2018kcc ; Khemchandani:2011et ; Doring:2008sv . Instead, we focus on the production of hidden-strange molecular-type pentaquarks with the $K^{*}\Sigma$ component. This leads to two possible $J^{P}$ quantum numbers for S-wave $K^{*}\Sigma$ molecular systems: $J^{P}=1/2^{-}$ and $3/2^{-}$. In this study, we designate two S-wave $K^{*}\Sigma$ molecular-type pentaquarks with $J^{P}=1/2^{-}$ and $3/2^{-}$ as $P_{s\bar{s}}[1/2^{-}]$ and $P_{s\bar{s}}[3/2^{-}]$, respectively.
In this work, we investigate the production of two possible hidden-strange molecular-type pentaquarks through the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction. We notice several previous experimental results, where the measured scattering cross section data for the $\pi^{-}p\rightarrow K^{*}\Sigma$ process were given Crennell:1972km ; CERN-CollegedeFrance-Madrid-Stockholm:1980mch ; CERN-CollegedeFrance-Madrid-Stockholm:1980ysu ; Dahl:1967pg ; Miller:1965 ; Abramovich:1972rq . It provides a good chance to combining with these experimental data to carry out the present research.
@@
For quantitatively calculating the $\pi^{-}p\rightarrow K^{*}\Sigma$ scattering process, we adopt the effective Lagrangian approach, with the aim of investigating the production of hidden-strange molecular pentaquarks $P_{s\bar{s}}$. Additionally, we will analyze the $\pi^{-}p\rightarrow K^{*}\Sigma\rightarrow K^{+}\pi^{-}\Sigma^{0}$ Dalitz process and assess its experimental feasibility, thereby offering valuable theoretical insights for future experimental studies.
@@
This paper is organized as follows. After the introduction, the Lagrangians and amplitudes used in this work are presented in Section II. The numerical results for the cross sections and the constituent counting rule are provided in Section III. The Dalitz process and its experimental feasibility are discussed in Section IV, followed by a brief summary in Section V.
II THE PRODUCTION OF THE $P_{s\bar{s}}$ VIA THE $\pi^{-}p\rightarrow K^{*}\Sigma$ REACTION
@@
The schematic tree-level Feynman diagrams for the $\pi^{-}p\rightarrow K^{*}\Sigma$ reaction are depicted in Fig. 1. These include $s$-channel $P_{s\bar{s}}$ ($\equiv P^{*}$) exchange, $u$-channel $\Sigma$ exchange, and $t$-channel $K$ and $K^{*}$ exchanges.
@@
To evaluate the pion-induced production of $K^{*}\Sigma$, the relevant Lagrangians are required, as used in previous studies Xiang:2020phx ; Kim:2015ita ; He:2017aps ; Xie:2014kja ; Ben:2023uev ; Wang:2019dsi ; Cheng:2016ddp . In general, similar to the case of the $P_{c}$ pentaquark state, the $J^{P}$ quantum numbers of the $P_{s\bar{s}}$ state can be either $\frac{3}{2}^{-}$ or $\frac{1}{2}^{-}$. Therefore, for the $s$-channel, both quantum number assignments for the $P_{s\bar{s}}$ should be considered in our calculations. These Lagrangians include
$\mathcal{L}^{3/2^{-}}_{K^{*}\Sigma P^{*}}$ $=$ $g_{K^{*}\Sigma P^{*}}^{3/2^{-}}\bar{P^{*}}_{\mu}{\Sigma}K^{*\mu}+\mathrm{H.c.},$ (1)
$\mathcal{L}^{3/2^{-}}_{\pi NP^{*}}$ $=$ $\frac{g_{\pi NP^{*}}^{3/2^{-}}}{m_{\pi}}\bar{N}\gamma_{5}\bm{\tau}\cdot\partial^{\mu}\bm{\pi}P^{*}_{\mu}+\mathrm{H.c.},$ (2)
$\mathcal{L}^{1/2^{-}}_{K^{*}\Sigma P^{*}}$ $=$ $g_{K^{*}\Sigma P^{*}}^{1/2^{-}}\bar{P^{*}}\gamma_{5}\gamma^{\mu}K^{*}_{\mu}\Sigma+\mathrm{H.c.},$ (3)
$\mathcal{L}^{1/2^{-}}_{\pi NP^{*}}$ $=$ $g_{\pi NP^{*}}^{1/2^{-}}\bar{N}\bm{\tau}\cdot\bm{\pi}P^{*}+\mathrm{H.c.}.$ (4)
@@
where $\bm{\tau}$ denotes the Pauli matrices, and the matrix $\bm{T}$ is defined as in Ref. Matsuyama:2006rp . The symbols $\bm{\pi}$, $K$, and $K^{*}$ represent the $\pi$, $K$, and $K^{*}$ meson fields, respectively. Similarly, $N$, $\Sigma$, and $P^{*}$ correspond to the nucleon, $\Sigma$ baryon, and $P_{s\bar{s}}$ fields, respectively.
Assuming the $P_{s\bar{s}}[3/2^{-}]$ is a pure $S$-wave molecular state of $K^{*}$ and $\Sigma$, the coupling constant $g_{K^{*}\Sigma P^{*}}^{3/2^{-}}$ can be estimated model-independently using the Weinberg compositeness criterion, yielding Weinberg:1965zz ; Baru:2003qq ; Lin:2017mtz ; Lin:2018kcc
$g_{K^{*}\Sigma P^{*}}^{2}$ $=$ $\frac{4\pi}{4M_{P^{*}}M_{\Sigma}}\frac{(M_{K^{*}}+M_{\Sigma})^{5/2}}{(M_{K^{*}}M_{\Sigma})^{1/2}}\sqrt{32\epsilon},$ (11)
$\epsilon$$\equiv$ $M_{K^{*}}+M_{\Sigma}-M_{P^{*}},$ (12)
where $M_{P^{*}}$, $M_{K^{*}}$, and $M_{\Sigma}$ represent the masses of the $P_{s\bar{s}}[3/2^{-}]$, $K^{*}$, and $\Sigma$, respectively, while $\epsilon$ denotes the $K^{*}\Sigma$ binding energy. Assuming the physical state is a pure $S$-wave hadronic molecule, the relative uncertainty of the above approximation for the coupling constant is given by $\sqrt{2\mu\epsilon}\cdot r$, where $\mu=\frac{m_{1}m_{2}}{m_{1}+m_{2}}$ is the reduced mass of the bound particles, and $r$ is the range of forces, which can be estimated as the inverse of the mass of the exchanged particle. For the $K^{*}\Sigma$ system, $r$ can be estimated as $1/m_{K}$ Lin:2018kcc .
Following Ref. Lin:2018kcc , we adopt the mass of the $P_{s\bar{s}}[3/2^{-}]$ as $M_{P^{*}}=2080\text{ MeV}$. Substituting this into Eq. (11), one obtains:
$g_{K^{*}\Sigma P^{*}}^{3/2^{-}}=1.72.$ (13)
Since the $P_{s\bar{s}}[1/2^{-}]$ can also be considered a $S$-wave molecular state of $K^{*}$ and $\Sigma$, we approximate the value of $g_{K^{*}\Sigma P^{*}}^{1/2^{-}}$ as 1.72. In the absence of experimental measurements for the partial width of the $P_{s\bar{s}}$ decays to $\pi N$, we treat the coupling constants $g_{\pi NP^{*}}^{3/2^{-}}$ and $g_{\pi NP^{*}}^{1/2^{-}}$ as free parameters to be determined through fitting.
@@
$F_{t/u}(q)=\left(\frac{\Lambda{{}^{2}_{t/u}}-m^{2}}{\Lambda{{}^{2}_{t/u}}-q^{2}}\right)^{2}.$ (16)
@@
$\mathcal{A}^{P^{*({3/2}^{-})}}_{s,\mu}$ $=$ $\sqrt{2}g_{K^{*}\Sigma P^{*}}^{{3/2}^{-}}\frac{g_{\pi NP^{*}}^{{3/2}^{-}}}{m_{\pi}}F_{s}(q)\gamma_{5}k_{1}^{\nu}$ (19)
×$\times\frac{(q\mkern-8.0mu/_{s}+m_{P^{*}})}{s-m^{2}_{P^{*}}+im_{P^{*}}\Gamma_{P^{*}}}\Delta_{\mu\nu},$
$\mathcal{A}^{P^{*({1/2}^{-})}}_{s,\mu}$ $=$ $-i\sqrt{2}g_{K^{*}\Sigma P^{*}}^{{1/2}^{-}}g_{\pi NP^{*}}^{{1/2}^{-}}F_{s}(q)\gamma_{5}\gamma_{\mu}$ (20)
×$\times\frac{(q\mkern-8.0mu/_{s}+m_{P^{*}})}{s-m^{2}_{P^{*}}+im_{P^{*}}\Gamma_{P^{*}}},$
$\mathcal{A}^{\Sigma}_{u,\mu}$ $=$ $-i\sqrt{2}\frac{g_{\Sigma\Sigma\pi}}{m_{\pi}}g_{K^{*}N\Sigma}F_{u}(q)\gamma^{5}\gamma^{\nu}k_{1\nu}\frac{(q\mkern-8.0mu/_{\Sigma}+m_{\Sigma})}{{u-m_{\Sigma}^{2}}}$ (21)
@@
$\mathcal{A}^{K}_{t,\mu}$ $=$ $i\sqrt{2}g_{K^{*}K\pi}g_{K\Sigma N}F_{t}(q)\frac{1}{{t-m_{K}^{2}}}\gamma_{5}$ (22)
@@
$\mathcal{A}^{K^{*}}_{t,\mu}$ $=$ $i\sqrt{2}g_{\pi K^{*}K^{*}}g_{K^{*}N\Sigma}F_{t}(q)(\gamma_{\xi}-\frac{k_{K^{*}{\Sigma}N}}{2m_{N}}\gamma_{\xi}q\mkern-8.0mu/_{K^{*}})$ (23)